OpenStudy (anonymous):

In right triangle DEF, d = 8, e = 10, f = 6. Which is cos D + cos F? 2.40 2.05 1.40 0.76 .

OpenStudy (anonymous):

D is the angle opposit d and F is the angle opposite f. e being the longest side is the hypotenuse. Now to calculate cosD we find adjacent/hypotenuse=6/10 and to calculate cosF we find adjacent/hypotenuse=8/10 so, cosD +cosF=0.6+.8=1.4

OpenStudy (anonymous):

thank you !